How much 60% disinfectant solution must be mixed with a 80% disinfectant solution to produce 25 gallons of a 72% disinfectant solution?
Answer by stanbon(75887)

How much 60% disinfectant solution must be mixed with a 80% disinfectant solution to produce 25 gallons of a 72% disinfectant solution?
-------
Equation:
active + active = active
0.60x + 0.80(25-x) = 0.72*25
------
Multiply thru by 100 to get:
60x + 80*25 - 80x = 72*25
-20x = -8*25
x = 10 gallons (amt. of 60% solution needed)
======================================
Cheers,
Stan H.
==============